8. Needham Pharmaceuticals has a profit margin of 3.5% and an equity multiplier of 1.6. Its sales are $110 million and it has total assets of $52 million. What is its Return on Equity (ROE)? Round your answer to two decimal places.

Ans: ___________%
Answer #1

total asset turnover=Sales/total asset

=(110/52)=2.11538462

ROE=profit margin*total asset turnover*equity multiplier

=3.5*2.11538462*1.6

=11.85%(Approx).